E. The doctor writes a prescription for an insufflation for 500 mg of Amoxicillin for an adult. The active ingredient Amoxicillin is combined with a medium of Calcium Carbonate. There is a 1:3 ratio of Amoxicillin to Calcium Carbonate. The same insufflation needs to be created for an infant; the total weight of the new insuffulation is 300 mg. What is weight of Calcium Carbonate for the new insuffulation and what is the weight of Amoxicillin for the new insuffulation?
1
Expert's answer
2012-09-13T11:49:46-0400
Let the weight of Amoxicillin for the new insuffulation is Wa and the
weight
of Calcium Carbonate is Wc.
Wc = 3 * Wa.
The total weight Wt = Wa + Wc is
300 mg.
Wa + Wc = 300
Wa + 3 * Wa = 300
4 * Wa = 300
Wa = 75
mg
Wc = 3 * Wa = 225 mg